OpenStudy (anonymous):

using identities simplify : (sin^2x-4cos^2x)/1-5cos^2x

OpenStudy (anonymous):

@Hero @satellite73

OpenStudy (anonymous):

dangit no one wants to try this one lol @lgbasallote

OpenStudy (lgbasallote):

how is the question? \[\huge \frac{\sin^2 x - 4\cos ^2 x}{1 - 5\cos^2 x}\] ???

OpenStudy (anonymous):

looks correct

OpenStudy (lgbasallote):

change sin^2 x into cos^2 x first \[\huge \implies \frac{ 1 - \cos^2 x - 4\cos^2 x}{1 - 5\cos^2 x}\] combie like terms \[\huge \implies \frac{1 - 5\cos^2 x}{1 - 5\cos^2 x}\] simplify \[\huge \implies \frac{\cancel{1 - 5\cos^2 x}}{\cancel{1 - 5\cos^2 x}} \implies 1\] does that help?

OpenStudy (anonymous):

oh your good lol

OpenStudy (lgbasallote):

heh i bet you knew this all along